OpenStudy (anonymous):

Use the properties of exponents to explain why 8 raised to the power of 1 over 3 is called the cube root of 8

OpenStudy (asnaseer):

HINT: what is \(x^{1/3}\times x^{1/3}\times x^{1/3}=?\)

OpenStudy (anonymous):

To be honest, I have no idea....

OpenStudy (asnaseer):

what laws of exponents do you know?

OpenStudy (anonymous):

Absolutely none, I hate just getting the answers but I have so much other work to do and this is the only thing I am stuck on.

OpenStudy (asnaseer):

I suggest you brush up on this first and then attempt this question. This site might prove useful: http://www.mathsisfun.com/exponent.html

OpenStudy (anonymous):

I don't have time, it's a test & I'm timed.

OpenStudy (confluxepic):

It is much better to learn by yourself and remember it instead of getting answers without explanation.

OpenStudy (anonymous):

Alright, lol. Try explaining it then.

OpenStudy (confluxepic):

You see the power of 3 is called cubed.

OpenStudy (confluxepic):

The power of 1/3 will be the cubed root.

OpenStudy (confluxepic):

It is just like x^2 and x^1/2.

OpenStudy (confluxepic):

x^1/2 is the same as the square root of x.

OpenStudy (anonymous):

I'm so confused

OpenStudy (confluxepic):

Remember the word root.

OpenStudy (confluxepic):

You see 1/2 is less than 1.

OpenStudy (confluxepic):

Instead of raising the value higher it makes the value lower.

OpenStudy (confluxepic):

Just like when you do a square root.

OpenStudy (confluxepic):

The number 3 can be associated with cubed when it comes to exponents.

OpenStudy (confluxepic):

It can be confusing.